# A Closer Look at Multimodal Representation Collapse

### v1

@@ 3.1 Polysemanticity and Cross-Modal Entanglements
We begin by showing that as the number of modalities increase, the proportion of cross-modal polysemantic neurons, i.e., those that encode features from more than one modality (as opposed to monosemantic neurons which encode exactly one feature from one modality) also increases (Lemma 1). This makes it difficult for the fusion head to independently control the contribution from a given modality without potential destructive interference from others (Theorem 1).
@@ 3.2 Rank Bottleneck
Lemma 2.
(Gradient Rank) The rank of gradient updates across iterations of SGD at layer $l$ is a convergent sequence with the following limit:
@@
Theorem 2 implies that since the AGOP is known to be the low-rank subspace that $W$ converges to under SGD (Radhakrishnan et al., 2024), the small distance (tighter bound) between the AGOP and ${\mathrm{\mathbf{w}}}$ implies that $W$ in fact converges to the polysemantic subspace ${\mathrm{\mathbf{w}}}$. In other words, SGD is more likely to parameterize $W$ with the low-rank polysemantic neurons than with high-rank monosemantic ones. Below we explore ways of breaking this implicit rank bottleneck to circumvent cross-modal polysemantic interference.
@@ Appendix B Proofs
The number of ways any two features can be selected from by $\varphi$ from $X$ such that both belong to different modalities is ($\geq{m\choose 2}(\dim f_{\text{min}})^{2}$, since there are (${m\choose 2}$ ways of choosing modality-pairs, and there are $\geq(\dim f_{\text{min}})^{2}$ ways of choosing feature pairs in each such combination. Now, it is these pairs of features that lead to cross-modal polysemantic collisions (through neuron subspaces ${\mathrm{\mathbf{w}}}_{p}$) during fusion in $\varphi$. Let the ambient dimension of the input to $\varphi$ be $F_{\text{dim}}={\displaystyle\sum_{i=1}^{m}\dim f_{i}}$. Then, for a given depth and width, the probability that a polysemantic weight subspace would represent features from two different modalities would be:
@@
Since noisy features are closer to random, they can get entangled with neurons representing predictive features from any modality if the corresponding neuron allows features up to $(1-K)$ units of deviations, according to the Johnson–Lindenstrauss lemma (Elhage et al., 2022), i.e., satisfying the following:
@@
With the elimination of noisy features, such entanglements following from superposition become less likely among predictive features across modalities, since they would normally require dedicated dimensions with strong deviations from perfect orthogonality, i.e., monosemantic neurons. This completes the proof of the lemma. ∎
Theorem 1 (Interference) As the number of cross-modal polysemantic collisions increase, the fraction of predictive conjugate features contributing to the reduction of the task loss decreases, resulting in the following limit:
@@ B.1 Rank Bottleneck
Theorem 2 (Polysemantic Bottleneck) Let $W$ be the weight matrix at a given layer of $\varphi$, and ${\mathrm{\mathbf{w}}}\leq W$ be any subspace in $W$. When the reduction in conditional cross-entropy $H({\mathrm{\mathbf{x}}};{\mathrm{\mathbf{y}}}|{\mathrm{\mathbf{z}}})$ provided (amount of unique label information held) by each feature is the same, i.e., $I({\mathrm{\mathbf{x}}};{\mathrm{\mathbf{y}}}|{\mathrm{\mathbf{z}}}_{1})=I({\mathrm{\mathbf{x}}};{\mathrm{\mathbf{y}}}|{\mathrm{\mathbf{z}}}_{2})=...=I({\mathrm{\mathbf{x}}};{\mathrm{\mathbf{y}}}|{\mathrm{\mathbf{z}}}_{k})$, at any iteration $n$ of SGD, the norm of the difference between ${\mathrm{\mathbf{w}}}$ and the average gradient outer product (AGOP) of the complete weight matrix $W$ is bounded as follows:
@@ B.2 Knowledge Distillation Frees Up Rank Bottlenecks
Theorem 3 (Dynamic Convergence Bound]). When the inputs to $\varphi$ are dynamic (for instance, when the unimodal representations are aligned via cross-modal knowledge distillation) under some distance metric $d$, then at any iteration $n$ of SGD, the norm of the difference between ${\mathrm{\mathbf{w}}}$ and the AGOP of $W$ is bounded as follows for all modalities $i,j\in M$ and datapoints ${\mathrm{\mathbf{x}}}\in X$:
@@
Every modality consists of both noisy and predictive features. If fusion collapses to a specific modality (target), it means that the modality contains more predictive information and less noise than the rest. Knowledge distillation to align the representations of the other modalities with the target would thus denoise the other modalities, allocating a larger fraction of the feature space of the modality-specific encodings of such modalities to predictive features. Since noisy features are closer to random, they can get entangled with predictive features from any modality if the corresponding neuron has a slight deviation from perfect orthogonality, according to the Johnson–Lindenstrauss lemma (Elhage et al., 2022). With the elimination of noisy features, such entanglements following from superposition become less likely among predictive features across modalities, since they would normally require dedicated dimensions with strong deviations from perfect orthogonality, i.e., monosemantic neurons. Therefore, as the unimodal representations get closer to each other through the implicit denoising mechanism of knowledge distillation, SGD gets increasingly compelled to parameterize the fusion head in a monosemantic manner.
As the representations from different modalities get closer to each other under the distance metric $d$, which is what effectively happens during cross-modal knowledge distillation, the proportion of monosemantic neurons in $W$ increases. This results in the AGOP of $W$ diverging away from its polysemantic subspaces ${\mathrm{\mathbf{w}}}$. In other words, knowledge distillation implicitly disentangles the cross-modal interferences by freeing rank bottlenecks and encouraging necessary monosemanticity, allowing for independent, modality-wise denoising of features along novel dimensions.
In other words, when cross-modal polysemantic weight matrices are rank-bottlenecked, the only solution to minimize the loss further is to allocate the noisy features new, independent dimensions in the latent space. However, this causes an increase in representation rank. To counteract this, knowledge distillation frees up the rank bottlenecks by down-weighting the rank-regularization term.
@@
An intuition for this can be developed based on the low-rank simplicity bias. A weighted general case of the Depth-Rank Duality result by (Sreelatha et al., 2024) follows as a consequence of cross-modal interferences, which can be written as:
$\underbrace{\min_{\phi,f}\operatorname{\mathcal{L}}(\varphi(X),Y)}_{\text{ERM}}+\left(1-\kappa\right)\alpha\min_{\phi}\sum_{d}\operatorname{rank}(\varphi_{l}),$
@@
where $\kappa$ is the strength of cross-modal knowledge distillation. Now, when cross-modal polysemantic interference happens despite there being extra available dimensions in the ambient space, it indicates a faulty capacity allocation (Scherlis et al., 2022) of the corresponding neurons by SGD, which follows from the Johnson–Lindenstrauss lemma (Elhage et al., 2022). It happens because SGD, by default, performs the aforementioned implicit weighted rank-regularization aside from ERM, with equal weights between ERM and the rank regularization term.
Knowledge distillation reduces the weight on the rank-regularization term, freeing up other dimensions for exploration, potentially containing higher rank solutions with more modalities. The rank de-regularization happens due to knowledge distillation having to denoise the student modality in order to align its representation with that of the teacher (Bishop, 1995). Once the optimizer converges to the modality-collapsed solution at the saddle point, the value of the loss and the rank of the solution balance each other out. Now, to explore new dimensions from the saddle point, the optimizer has to explore a novel direction in the parameter space, which leads to a reduction in loss but a simultaneous increase in rank. The reason behind this saddle-geometry is the presence of noisy features from the one modality in entanglement with the predictive features from the another, which results in an adversarial minimax game between the two. On either side of the saddle point along the unexplored dimensions are predictive features of the former modality which could potentially minimize the task loss, but is not taken into account due to rank regularization.
Since the rank-regularization happens when ${d(\tilde{{\mathrm{\mathbf{x}}}}_{i},\tilde{{\mathrm{\mathbf{x}}}}_{j})\to\epsilon}$, it implies that knowledge distillation down-weights this regularization by disentanglement and denoising of the bases of the modalities from which the noise features originate.
@@
Specifically, under a dynamic input space approaching a bounded neighborhood $\epsilon$, the number of features encoded in any polysemantic subspace gets bounded by a constant (say, $k_{\mathrm{\mathbf{w}}}$) as the $n\to\infty$. So, the RHS in Theorem 2 can be rewritten as:
@@
where $\kappa=k_{\mathrm{\mathbf{w}}}/\dim({\mathrm{\mathbf{w}}})$ is a constant is both $k_{\mathrm{\mathbf{w}}}$ and $\dim({\mathrm{\mathbf{w}}})$ are constants. Finally, following from the above, when ${d(\tilde{{\mathrm{\mathbf{x}}}}_{i},\tilde{{\mathrm{\mathbf{x}}}}_{j})\to\epsilon}$, Equation 4 from the proof of Theorem 2 can be expressed in terms of $\kappa$ as:

### v2

@@ 3.1 Polysemanticity and Cross-Modal Entanglements
We begin by showing that as the number of modalities increase, the proportion of cross-modal polysemantic neurons, i.e., those that encode features from more than one modality (as opposed to monosemantic neurons which encode exactly one feature from one modality) also increases (Lemma 1). This makes it difficult for the fusion head to independently control the contribution from a given modality without potential destructive interference from others (Theorem 1). It is important to note that the results presented presuppose the occurrence of polysemanticity, i.e., the number of task-relevant features in $X$ is greater than the number of neurons in any layer, something that is most often known to hold in practice (Scherlis et al., 2022).
@@
Since we are interested in the fraction of neurons that are cross-modal polysemantic in the space of all polysemantic neurons, $p(\cdot)$ is defined over the space of polysemantic neurons only. Also, Lemma 1 deals specifically with 2-semantic neurons, i.e., those that simultaneously encode 2 features, which is the most likely form of polysemanticity in the combinatorial space of cross-modal polysemanticities for any value of $m$.
@@ 3.2 Rank Bottleneck
Lemma 2 (Gradient Rank).
The rank of gradient updates across iterations of SGD at layer $l$ is a convergent sequence with the following limit:
@@
Theorem 2 implies that since the AGOP is known to be the low-rank subspace that $W$ converges to under SGD (Radhakrishnan et al., 2024), the small distance (tighter bound) between the AGOP and subspaces $\mathrm{\mathbf{w}}$} with higher $\gamma({\mathrm{\mathbf{w}}})$ implies that $W$ in fact converges to those subspaces $\mathrm{\mathbf{w}}$} with higher degrees of polysemanticity $\gamma({\mathrm{\mathbf{w}}})$. In other words, SGD is more likely to parameterize $W$ with the low-rank polysemantic neurons than with high-rank monosemantic ones. The implication of this is that, if we specifically consider the noisy-predictive type of cross-modal polysemantic neurons, they will be the first to get eliminated among all cross-modal polysemantic neurons due to the low-rank simplicity bias, as they either do not contribute to loss reduction, or do so negatively. Below we explore ways of breaking this implicit rank bottleneck to circumvent cross-modal polysemantic interference among noisy and predictive features.
@@ Appendix B Proofs
The number of ways any two features can be selected by $\varphi$ from $X$ such that both belong to different modalities is $\geq{m\choose 2}(\dim f_{\text{min}})^{2}$, since there are $m\choose 2$} ways of choosing modality-pairs, and there are $\geq(\dim f_{\text{min}})^{2}$ ways of choosing feature pairs in each such combination. Now, it is these pairs of features that lead to cross-modal polysemantic collisions (through neuron subspaces ${\mathrm{\mathbf{w}}}_{p}$) during fusion in $\varphi$. Let the ambient dimension of the input to $\varphi$ be $F_{\text{dim}}={\displaystyle\sum_{i=1}^{m}\dim f_{i}}$. Then, for a given depth and width, the probability that a polysemantic weight subspace would represent features from two different modalities would be:
@@
Since noisy features are closer to random, they can get entangled with neurons representing predictive features from any modality if the corresponding neuron allows features up to (1K)(1-K) units of deviations, according to the Johnson–Lindenstrauss lemma (Elhage et al., 2022), i.e., satisfying the following:
@@
With the elimination of noisy features, such entanglements following from superposition become less likely among predictive features across modalities, since they would normally require dedicated dimensions with strong orthogonality, i.e., monosemantic neurons. This completes the proof of the lemma. ∎
Theorem 1 (Interference). As the number of cross-modal polysemantic collisions increase, the fraction of predictive conjugate features contributing to the reduction of the task loss decreases, resulting in the following limit:
@@ B.1 Rank Bottleneck
Theorem 2 (Polysemantic Bottleneck). Let $W$ be the weight matrix at a given layer of $\varphi$, and ${\mathrm{\mathbf{w}}}\leq W$ be any subspace in $W$. When the reduction in conditional cross-entropy $H({\mathrm{\mathbf{x}}};{\mathrm{\mathbf{y}}}|{\mathrm{\mathbf{z}}})$ provided (amount of unique label information held) by each feature is the same, i.e., $I({\mathrm{\mathbf{x}}};{\mathrm{\mathbf{y}}}|{\mathrm{\mathbf{z}}}_{1})=I({\mathrm{\mathbf{x}}};{\mathrm{\mathbf{y}}}|{\mathrm{\mathbf{z}}}_{2})=...=I({\mathrm{\mathbf{x}}};{\mathrm{\mathbf{y}}}|{\mathrm{\mathbf{z}}}_{k})$, at any iteration $n$ of SGD, the norm of the difference between $\mathrm{\mathbf{w}}$} and the average gradient outer product (AGOP) of the complete weight matrix $W$ is bounded as follows:
@@ B.2 Knowledge Distillation Frees Up Rank Bottlenecks
Theorem 3 (Dynamic Convergence Bound). When the inputs to $\varphi$ are dynamic (for instance, when the unimodal representations are aligned via cross-modal knowledge distillation) under some distance metric $d$, then at any iteration $n$ of SGD, the norm of the difference between $\mathrm{\mathbf{w}}$} and the AGOP of $W$ is bounded as follows for all modalities $i,j\in M$ and datapoints ${\mathrm{\mathbf{x}}}\in X$:
@@
A weighted general case of the Depth-Rank Duality result by (Sreelatha et al., 2024) follows as a consequence of cross-modal interferences, which can be written as:
$\underbrace{\min_{\varphi,f}\operatorname{\mathcal{L}}(\varphi(X),Y)}_{\text{ERM}}+\alpha\min_{\varphi}\sum_{l}\operatorname{rank}(\varphi),$
where $\alpha$ is a function measuring the compactness of the embedding space, in our case, through the pairwise distance between points, $d(\tilde{{\mathrm{\mathbf{x}}}}_{i},\tilde{{\mathrm{\mathbf{x}}}}_{j})$, computed via a given iterate of $\varphi$, i.e.,
$\alpha=h_{\varphi}(d(\tilde{{\mathrm{\mathbf{x}}}}_{i},\tilde{{\mathrm{\mathbf{x}}}}_{j}))$
@@
When cross-modal polysemantic interference happens despite there being extra available dimensions in the ambient space, it indicates a faulty capacity allocation (Scherlis et al., 2022) of the corresponding neurons by SGD, which follows from the Johnson–Lindenstrauss lemma (Elhage et al., 2022). It happens because SGD, by default, performs the aforementioned implicit weighted rank-regularization aside from ERM, with the highest possible value of $\alpha$ such that it encourages the smallest possible rank for a target ERM solution. However, if it is known a priori that the pairwise distances $d(\tilde{{\mathrm{\mathbf{x}}}}_{i},\tilde{{\mathrm{\mathbf{x}}}}_{j})$ approach some constant neighborhood $\epsilon$, we get:
$\alpha_{1}>\alpha_{2}>...>\alpha_{n}=h_{\varphi_{n}}(\epsilon)=A,$
where $A$ is a constant. Following from this, under the given limit, the rank regularization term in the optimization objective of Depth-Rank Duality gets relaxed as:
$\lim_{{d(\tilde{{\mathrm{\mathbf{x}}}}_{i},\tilde{{\mathrm{\mathbf{x}}}}_{j})\to\epsilon}}\alpha\min_{\varphi}\sum_{l}\operatorname{rank}(\varphi)=A\min_{\varphi}\sum_{l}\operatorname{rank}(\varphi),$
which ultimately implies that for any $\tilde{\varphi}$ inducing pairwise distances $d(\tilde{{\mathrm{\mathbf{x}}}}_{i},\tilde{{\mathrm{\mathbf{x}}}}_{j})>\epsilon$ will have $\operatorname{rank}(\tilde{\varphi})<\operatorname{rank}(\varphi)$. As representations from different modalities get closer to each other under the distance metric $d$, which is what effectively happens during cross-modal knowledge distillation, the increase in rank encourages a consequent increase in the proportion of monosemantic neurons in $W$. This results in the AGOP of $W$ diverging away from its polysemantic subspaces ${\mathrm{\mathbf{w}}}_{p}$, with the size of such polysemantic subspaces decreasing as follows:
$\left\lVert\tilde{{\mathrm{\mathbf{w}}}}_{p}-\sum_{{\mathrm{\mathbf{x}}}\in X}\nabla\tilde{\varphi}_{W}({\mathrm{\mathbf{x}}})\nabla\tilde{\varphi}_{W}({\mathrm{\mathbf{x}}})^{T}\right\rVert<\lim_{d(\tilde{{\mathrm{\mathbf{x}}}}_{i},\tilde{{\mathrm{\mathbf{x}}}}_{j})\to\epsilon}\left\lVert{\mathrm{\mathbf{w}}}_{p}-\sum_{{\mathrm{\mathbf{x}}}\in X}\nabla\varphi_{W}({\mathrm{\mathbf{x}}})\nabla\varphi_{W}({\mathrm{\mathbf{x}}})^{T}\right\rVert,$
Given the above constraint on the size of polysemantic bases in $W$ as $d(\tilde{{\mathrm{\mathbf{x}}}}_{i},\tilde{{\mathrm{\mathbf{x}}}}_{j})\to\epsilon$, the size of any feature subspace $\mathrm{\mathbf{w}}$}, i.e., the number of features that can be encoded by any $\mathrm{\mathbf{w}}$} approaches some constant upper bound corresponding to $\epsilon$:
$\lim_{{d(\tilde{{\mathrm{\mathbf{x}}}}_{i},\tilde{{\mathrm{\mathbf{x}}}}_{j})\to\epsilon}}|{\mathrm{\mathbf{w}}}\cap X|=|{\mathrm{\mathbf{w}}}_{\epsilon}\cap X|=k_{\mathrm{\mathbf{w}}},$
@@
where $|{\mathrm{\mathbf{w}}}_{\epsilon}|$ is the size of the feature subspace in the neigborhood $\epsilon$. Since both $|{\mathrm{\mathbf{w}}}_{\epsilon}|$ and $|X|$ are constants, $k_{\mathrm{\mathbf{w}}}$ is also a constant. It thus follows that, under a dynamic input space approaching a bounded neighborhood $\epsilon$, the number of features encoded in any polysemantic subspace also gets bounded by $k_{\mathrm{\mathbf{w}}}$ as $n\to\infty$. So, the RHS in Theorem 2 can be rewritten as:
@@
where $\kappa=k_{\mathrm{\mathbf{w}}}/\dim({\mathrm{\mathbf{w}}})$ is a constant as both $k_{\mathrm{\mathbf{w}}}$ and $\dim({\mathrm{\mathbf{w}}})$ are constants. Finally, following from the above, when $d(\tilde{{\mathrm{\mathbf{x}}}}_{i},\tilde{{\mathrm{\mathbf{x}}}}_{j})\to\epsilon$}, Equation 4 from the proof of Theorem 2 can be expressed in terms of $\kappa$ as:
@@ B.3 Additional Remarks
Intuition behind the Dynamic Convergence Bound: Every modality consists of both noisy and predictive features. If fusion collapses to a specific modality (target), it means that the modality contains more predictive information and less noise than the rest. Knowledge distillation to align the representations of the other modalities with the target would thus denoise the other modalities, allocating a larger fraction of the feature space of the modality-specific encodings of such modalities to predictive features. Since noisy features are closer to random, they can get entangled with predictive features from any modality if the corresponding neuron has a slight deviation from perfect orthogonality, according to the Johnson–Lindenstrauss lemma (Elhage et al., 2022). With the elimination of noisy features, such entanglements following from superposition become less likely among predictive features across modalities, since they would normally require dedicated dimensions with strong orthogonality, i.e., monosemantic neurons. Therefore, as the unimodal representations get closer to each other through the implicit denoising mechanism of knowledge distillation, SGD gets increasingly compelled to parameterize the fusion head in a monosemantic manner.
In other words, knowledge distillation implicitly disentangles the cross-modal interferences by freeing rank bottlenecks and encouraging necessary monosemanticity, allowing for independent, modality-wise denoising of features along novel dimensions.
When cross-modal polysemantic weight matrices are rank-bottlenecked, the only solution to minimize the loss further is to allocate the noisy features new, independent dimensions in the latent space. However, this causes an increase in representation rank. To counteract this, knowledge distillation frees up the rank bottlenecks by down-weighting the rank-regularization term.
Distillation Denoising Conjecture: Knowledge distillation reduces the weight on the rank-regularization term, freeing up other dimensions for exploration, potentially containing higher rank solutions with more modalities. The rank de-regularization happens due to knowledge distillation having to denoise the student modality in order to align its representation with that of the teacher (Bishop, 1995), as also indicated by our empirical observations in Section 4.3. Given this, we conjecture that knowledge distillation allows the representation of the noisy components of the teacher modality as a transformed version of the student noise, thereby eliminating the need for encoding noisy features from every modality in the neurons encoding the student modality. This can be formally stated as follows:
$\lim_{f_{s}(x_{s})\to f_{t}(x_{t})}f_{t}(x_{t}\hat{\eta})=g(x_{s}\hat{\eta});\;\;\forall x_{s}\in X_{s},x_{t}\in X_{t}$
where $X_{s}$ and $X_{t}$ are respectively the teacher and student modalities, $f_{s}$ and $f_{t}$ are respectively the teacher and student encoders, $\hat{\eta}$ represents the noisy components of a modality, and $g$ is the transformation function relating the student noise with the teacher noise embedding.
Loss Landscape Geometry: Once the optimizer converges to the modality-collapsed solution, the value of the loss and the rank of the solution balance each other out. Now, to explore further, the optimizer has to move along a novel direction in the parameter space which leads to a reduction in loss but a simultaneous increase in rank. The reason behind this saddle-geometry is the presence of noisy features from one modality in entanglement with the predictive features from another, which results in an adversarial minimax game between the two. On either side of the saddle point along the unexplored dimensions are predictive features of the former modality which could potentially minimize the task loss, but is not taken into account due to rank regularization. Since, according to Theorem 3, subspaces in $\varphi$ get decreasingly polysemantic when $d(\tilde{{\mathrm{\mathbf{x}}}}_{i},\tilde{{\mathrm{\mathbf{x}}}}_{j})\to\epsilon$}, it implies that knowledge distillation down-weights this regularization by disentanglement and denoising of the bases of the modalities from which the noisy features originate.
A Note on Figure 2: Under disentangled polysemanticity, although the noisy and predictive features may still be encoded via the same neuron, they map to different regions in the activation space of the neuron, leading to a feature-wise separable effect. For instance, the coefficients of the predictive features may have a relatively higher magnitude, implying that in order to activate the noisy component of the neuron, the degree of noise in the input would have to be much higher than its predictive counterpart.